OpenStudy (anonymous):

Integrating... (missing some kind of trig-related trick here) \[\int 5\cdot sin(8x)\cdot cos(5x)\cdot dx \] (medal for showing related substitutions :-) )

OpenStudy (goformit100):

∫5⋅sin(8x)⋅cos(5x)⋅dx

OpenStudy (turingtest):

a trig identity should solve your problem here\[\sin a\cdot\cos b=\frac12[\sin(a+b)+\sin(a-b)]\]

OpenStudy (turingtest):

for future reference, I keep this file handy at all times http://tutorial.math.lamar.edu/pdf/Trig_Cheat_Sheet.pdf

OpenStudy (anonymous):

Ah... Ty @TuringTest :-)